Hint Ch 29 > What is the slope of the given supply curve? 15 12, 515) 10 Ad (9, $9) 10 15 Quantity about us careers privacy policy terms of use contact us help
Answer #1

Slope of supply curve = (Y2 - Y1)/(X2 - X1)

Y1 = 9 , Y2 = 15, X1 = 9 and X2 = 12

Slope = (15 - 9)/(12 - 9) = 2